# On Discrete Morse-Bott Theory

### v1

@@ 4. Natural properties of discrete Morse-Bott functions
Let $f:\mathbb{K}\to\mathbb{R}$ be a function and $L$ a subcomplex of $\mathbb{K}$. Then $c^{L}_{k},m^{L}_{k},u^{L}_{k},d^{L}_{k}$ are defined as follows:
$c^{L}_{k}$ $:=\#\{\text{$k$-cells of }L\}$
$m^{L}_{k}$ $:=\#\{\text{critical $k$-cells for $f$ in }L\}$
$u^{L}_{k}$ $:=\sum_{\sigma^{k}\in L}\#\{\tau\in L\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}$
$d^{L}_{k}$ $:=\sum_{\sigma^{k}\in L}\#\{\nu\in L\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}$
Then the following relations hold:
$c^{L}_{k}$ $=m^{L}_{k}+u^{L}_{k}+d^{L}_{k}$
$u^{L}_{k}$ $=d^{L}_{k+1}$
@@
$c_{k}:=c^{\mathbb{K}}_{k},\quad m_{k}:=m^{\mathbb{K}}_{k},\quad u_{k}:=u^{\mathbb{K}}_{k},\quad d_{k}:=d^{\mathbb{K}}_{k}.$
@@
Let $f$ be a discrete Morse–Bott function on a finite CW complex $\mathbb{K}$. For any collection $C\in\mathcal{C}_{f}$, there is a polynomial $r(t)$ with nonnegative integer coefficients such that
$\sum_{k=0}^{\dim\mathbb{K}}\#\{\sigma\in C^{\mathrm{red}}\mid\dim\sigma=k\}t^{k}=P_{t}(C^{\mathrm{red}})+(1+t)r(t)$
@@
For any integer $k\in\mathbb{Z}_{>0}$, the following two short exact sequences hold:
$0\rightarrow Z^{C}_{k}\rightarrow C_{k}(C^{\mathrm{red}})\rightarrow B_{k-1}^{C}\rightarrow 0$
@@
Then we obtain the following equality for any integer $k\in\mathbb{Z}_{>0}$:
$\operatorname{rank}C_{k}(C^{\mathrm{red}})=\operatorname{rank}Z^{C}_{k}+\operatorname{rank}B_{k-1}^{C}$
@@
$\sum_{k=0}^{\dim\mathbb{K}}\#\{\sigma\in C^{\mathrm{red}}\mid\dim\sigma=k\}t^{k}-P_{t}(C^{\mathrm{red}})$ $=\sum_{k=0}^{\dim\mathbb{K}}\{\operatorname{rank}C_{k}(C^{\mathrm{red}})-b^{C}_{k}\}t^{k}$
@@
By $\operatorname{rank}B_{k-1}^{C}\geq 0$ for any integer $k\in\mathbb{Z}_{>0}$, any coefficient of $r(t)$ is non-negative. ∎
@@
where $P_{t}(\mathbb{K})=\sum_{k}b_{k}t^{k}$ is the Poincaré polynomisal of $\mathbb{K}$ and
$r(t)=\sum_{k=1}^{\dim\mathbb{K}}\bigl(\operatorname{rank}B_{k-1}-d_{k}\bigr).$
Although this proof is similar to that of [2004MorseIneq, Theorem 3.36] in the continuous setting, to the best of the authors’ knowledge, no direct proof has been given in the literature, so we include it here for completeness.
@@
For any $k\in\mathbb{Z}_{>0}$, from the two short exact sequences
@@
we obtain the following equality:
@@
$=(1+t)\sum_{k=1}^{\dim\mathbb{K}}(\operatorname{rank}B_{k-1}-d_{k})t^{k-1}$
@@
For a subcomplex $L$ of $\mathbb{K}$, for any $k\in\mathbb{Z}_{\geq 0}$, denote by $\partial_{L,k}:C_{k}(L)\to C_{k-1}(L)$ the boundary homomorphism. For any $k\in\mathbb{Z}_{\geq 0}$, since $\operatorname{rank}B_{k-1}-d_{k}=u_{k}+m_{k}-\operatorname{rank}Z_{k}$, it suffices to show the inequality
(2) $u_{k}+m_{k}-\operatorname{rank}Z_{k}\;\geq\;0$
@@
$f(\sigma)=\max\{f(\alpha)\mid\alpha\text{ is a maximal cell in }K\}.$
Then $K_{\sigma}:=K\setminus\{\sigma\}$ is a subcomplex of $\mathbb{K}$. Put $p:=\dim\sigma$. Since $B$ is a minimal subcomplex which dose not satisfy (1), the sub complex $K\setminus\{\sigma\}$ satisfies (1). Since $K_{\sigma}$ is obtained from $K$ by removing the single cell $\sigma$, we obtain that $\operatorname{rank}\ker\partial_{K,p}-\operatorname{rank}\ker\partial_{K_{\sigma},p}\in\{0,1\}$. Therefore, the inequalty (1) implies the following equality:
@@
$x:=\sum_{\alpha^{k}\in K}a_{\alpha}\alpha\in\ker\partial_{K,k},\qquad a_{\alpha}\in\mathbb{Z},$
@@
Let $f$ be a discrete Morse–Bott function on a finite CW complex $\mathbb{K}$ and $x\in Z_{k}\setminus\{0\}$. Write
@@
and $\alpha\in C^{\mathrm{red}}$ for some $C\in\mathcal{C}_{f}$, the sum
$x^{\mathrm{red}}:=\sum_{\sigma\in A\cap C^{\mathrm{red}}}a_{\sigma}\sigma$
satisfies $x^{\mathrm{red}}\in Z^{C}_{k}\setminus\{0\}$.
@@
First, we show the following statement.
@@
$\{\nu\prec\sigma\mid\nu\in C^{\mathrm{red}}\}=\emptyset$ for any cell $\sigma\in A\setminus C^{\mathrm{red}}$.
@@
Fix any cell $\sigma\in A\setminus C^{\mathrm{red}}$. Define $N_{\sigma}:=\{\nu\prec\sigma\mid\nu\in C^{\mathrm{red}}\}$. Assume that $N_{\sigma}\neq\emptyset$. Fix any cell $\nu\in N_{\sigma}$. Since $\nu\in C^{\mathrm{red}}$, the cell $\nu$ is not upward noncritical and so
@@
Because $\sigma\in A$, by the maximality (5) of $\alpha$, it follows $f(\nu)=f(\sigma)$ and so that $\sigma\in C\setminus C^{\mathrm{red}}$.
@@
$\{\nu\in\mathbb{K}\mid\sigma\in A,\;\nu\prec\sigma\}$ $=\{\nu\in\mathbb{K}\mid\sigma\in A,\;\nu\prec\sigma,\;\nu\notin C^{\mathrm{red}}\}$
$\quad\cup\{\nu\in\mathbb{K}\mid\sigma\in A\cap C^{\mathrm{red}},\;\nu\prec\sigma,\;\nu\in C^{\mathrm{red}}\}.$
@@
$=\sum_{\sigma\in A\cap C^{\mathrm{red}}}\sum_{\begin{subarray}{c}\nu\prec\sigma\\ \nu\in C^{\mathrm{red}}\end{subarray}}a_{\sigma}[\sigma:\nu]\nu+\sum_{\sigma\in A}\sum_{\begin{subarray}{c}\nu\prec\sigma\\ \nu\notin C^{\mathrm{red}}\end{subarray}}a_{\sigma}[\sigma:\nu]\nu$
@@
$0$ $=\sum_{\sigma\in A\cap C^{\mathrm{red}}}\sum_{\begin{subarray}{c}\nu\prec\sigma\\ \nu\in C^{\mathrm{red}}\end{subarray}}a_{\sigma}[\sigma:\nu]\nu=\partial^{C}_{k}x^{\mathrm{red}}$
@@
Let $f:\mathbb{K}\to\mathbb{R}$ be a function and $L$ a subcomplex of $\mathbb{K}$. Then $\tilde{m}^{L}_{k},\tilde{u}^{L}_{k},\tilde{d}^{L}_{k}$ are defined as follows:
$\tilde{m}^{L}_{k}$ $:=\sum_{C\in\mathcal{C}_{f}}\#\{\sigma^{k}\in L\mid\sigma\in C^{\mathrm{red}}\}$
$\tilde{u}^{L}_{k}$ $:=\sum_{\sigma^{k}\in L}\#\{\tau\in L\mid\sigma\overset{\mathrm{snc}}{\prec}\tau\}$
$\tilde{d}^{L}_{k}$ $:=\sum_{\sigma^{k}\in L}\#\{\nu\in L\mid\nu\overset{\mathrm{snc}}{\prec}\sigma\}$
Then the following relations hold:
$c^{L}_{k}$ $=\tilde{m}^{L}_{k}+\tilde{u}^{L}_{k}+\tilde{d}^{L}_{k}$
$\tilde{u}^{L}_{k}$ $=\tilde{d}^{L}_{k+1}$
@@
Let $f$ be a discrete Morse–Bott function on a finite CW complex $\mathbb{K}$. Then there is a polynomial $R(t)$ with nonnegative integer coefficients such that
(6) $\sum_{C\in\mathcal{C}_{f}}P_{t}(C^{\mathrm{red}})\;=\;P_{t}(\mathbb{K})+(1+t)R(t)$
@@
The relation $\nabla_{s}f=-\nabla g$ implies $m^{g}_{k}=\tilde{m}_{k}$. Therefore, the following holds:
@@
$\implies$ $\sum_{k=0}^{\dim\mathbb{K}}\sum_{C\in\mathcal{C}_{f}}\#\{\sigma\in C^{\mathrm{red}}\mid\dim\sigma=k\}t^{k}=P_{t}(\mathbb{K})+(1+t)r(t)$
$\implies$ $\sum_{C\in\mathcal{C}_{f}}(P_{t}(C^{\mathrm{red}})+(1+t)\sum_{k=1}^{\dim\mathbb{K}}(\operatorname{rank}B_{k-1}^{C})t^{k-1})=P_{t}(\mathbb{K})+(1+t)r(t)$
$\implies$ $\sum_{C\in\mathcal{C}_{f}}P_{t}(C^{\mathrm{red}})=P_{t}(\mathbb{K})+(1+t)$
$\hskip 110.0pt\sum_{k=1}^{\dim\mathbb{K}}(\operatorname{rank}B_{k-1}-d^{g}_{k}-\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}B_{k-1}^{C})t^{k-1}$
@@
$\operatorname{rank}C_{k}(C^{\mathrm{red}})$ $=\operatorname{rank}Z^{C}_{k}+\operatorname{rank}B_{k-1}^{C}$
$c^{L}_{k}$ $=\tilde{m}^{L}_{k}+\tilde{u}^{L}_{k}+\tilde{d}^{L}_{k}$
@@
$\operatorname{rank}B_{k-1}-\tilde{d}_{k}-\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}B_{k-1}^{C}=\tilde{u}_{k}+\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}Z^{C}_{k}-\operatorname{rank}Z_{k}$
For any subcomplex $L$ of $\mathbb{K}$, put $\mathcal{C}_{L}:=\mathcal{C}_{f\mid_{L}}$. Then it suffices to show the following inequality:
(7) $\tilde{u}^{L}_{k}+\sum_{C\in\mathcal{C}_{L}}\operatorname{rank}Z^{C}_{k}-\operatorname{rank}\ker\partial_{L,k}\geq 0$
@@
$M:=\{\sigma\in K\mid\sigma\text{ is maximal in }K\}.$
@@
Let $C\in\mathcal{C}_{f}$ be a collection with $C^{\mathrm{red}}=\{\sigma,\tau\}$ with $\sigma^{p}\overset{\mathrm{reg}}{\prec}\tau$. By definition of regularity of faces,
we obtain that $P_{t}(C^{\mathrm{red}})=0$.
Suppose that $f$ is Morse.
@@
By Theorem 3.2, for any collection $C\in\mathcal{C}$, we obtain that either #C=1\#C=1 or both #C=2\#C=2 and $C=C^{\mathrm{red}}$. For any collection $C\in\mathcal{C}$ with #C=2\#C=2, we have that $P_{t}(C^{\mathrm{red}})=0$. For any collection $C\in\mathcal{C}$ with #C=1\#C=1 and $C^{\mathrm{red}}=\emptyset$, we obtain that $P_{t}(C^{\mathrm{red}})=0$. By Remark 3, any collection $C\in\mathcal{C}$ consists of one critical cell if and only if $C=\#C^{\mathrm{red}}=1$. By definition of $m_{k}$, the desired equality holds. ∎
@@
For any collection $C\in\mathcal{C}_{f}$ with #C=1\#C=1, the definition of $B_{k-1}^{C}$ implies $\operatorname{rank}B_{k-1}^{C}=0$. Therefore, we have the following equalty:
$\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}B_{k-1}^{C}=\sum_{C\in\mathcal{C}_{f},\#C=2}\operatorname{rank}B_{k-1}^{C}$
On the other hand, the following equality holds:
$\sum_{\sigma\in\mathbb{K}}\#\{\nu\in\mathbb{K}\mid f(\nu)=f(\sigma),\nu\prec\sigma\}$
$=$ $\sum_{C\in\mathcal{C}}\sum_{\sigma\in C}\#\{\nu\in\mathbb{K}\mid f(\nu)=f(\sigma),\nu\prec\sigma\}$
$=$ $\sum_{C\in\mathcal{C}}\sum_{\sigma\in C}\#\{\nu\in C\mid\nu\prec\sigma\}$
$=$ $\sum_{C\in\mathcal{C}}\sum_{\sigma\in C^{\mathrm{red}}}\#\{\nu\in C^{\mathrm{red}}\mid\nu\prec\sigma\}$
In the sequence of equalities above, the fact that any collection is contained in the inverse image of a point is applied in the second line. Since $\operatorname{rank}B_{k-1}^{C}=0$ for any collection $C\in\mathcal{C}$ with #C=1\#C=1, the third equality holds.
It suffices to show that $\operatorname{rank}B_{k-1}^{C}=\sum_{\sigma^{k}\in C^{\mathrm{red}}}\#\{\nu\in C^{\mathrm{red}}\mid\nu\prec\sigma\}$ for any collection $C\in\mathcal{C}_{f}$ with #C=2\#C=2.
Fix any collection $C\in\mathcal{C}_{f}$ with #C=2\#C=2. If any of $\operatorname{rank}B_{k-1}^{C}$ and $\sum_{\sigma^{k}\in C^{\mathrm{red}}}\#\{\nu\in C^{\mathrm{red}}\mid\nu\prec\sigma\}$) is non-zero, then it must contains a k-cell and a (k-1)-cell in $C$. By #C=2\#C=2, both $\operatorname{rank}B_{k-1}^{C}$ and $\sum_{\sigma^{k}\in C^{\mathrm{red}}}\#\{\nu\in C^{\mathrm{red}}\mid\nu\prec\sigma\}$ are either zero or one.
Suppose that $\sum_{\sigma^{k}\in C^{\mathrm{red}}}\#\{\nu\in C^{\mathrm{red}}\mid\nu\prec\sigma\}$ is one. Then there are cells $\nu\overset{\mathrm{nc}}{\prec}\sigma\in C$. By definition of regularity, we have $\nu\overset{\mathrm{reg}}{\prec}\sigma\in C$. Then $\partial^{C}\sigma=a_{\nu}\nu$ for some non-zero $a_{\nu}$. This implies $\operatorname{rank}B_{k-1}^{C}=1$.
Conversely, suppose that $\operatorname{rank}B_{k-1}^{C}=1$. By definition of $\partial^{C}$, there are cells $\nu\prec\sigma^{k}\in C=C^{\mathrm{red}}$. This means that $\sum_{\sigma^{k}\in C^{\mathrm{red}}}\#\{\nu\in C^{\mathrm{red}}\mid\nu\prec\sigma\}=1$.

### v2

@@ 4. Natural properties of discrete Morse-Bott functions
Let $f:\mathbb{K}\to\mathbb{R}$ be a function, $k\in\{0,1,\dots,\dim\mathbb{K}\}$ an integer, and $J$ a subcomplex of $\mathbb{K}$. Then $c^{J}_{k},m^{J}_{k},u^{J}_{k},d^{J}_{k}$ are defined as follows:
$c^{J}_{k}$ $:=\#\{\text{$k$-cells of }J\}$
$m^{J}_{k}$ $:=\#\{\text{critical $k$-cells for $f|_{J}$}\}$
$u^{J}_{k}$ $:=\sum_{\sigma^{k}\in J}\#\{\tau\in J\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}$
$d^{J}_{k}$ $:=\sum_{\sigma^{k}\in J}\#\{\nu\in J\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}$
If $f$ is Morse, then the following relations hold:
$c^{J}_{k}$ $=m^{J}_{k}+u^{J}_{k}+d^{J}_{k}\hskip 20.0pt\text{for any $k\in\{0,1,\dots,\dim\mathbb{K}\}$}$
$u^{J}_{k}$ $=d^{J}_{k+1}\hskip 61.0pt\text{for any $k\in\{0,1,\dots,\dim\mathbb{K}-1\}$}$
@@
$c_{k}:=c^{\mathbb{K}}_{k},\quad m_{k}:=m^{\mathbb{K}}_{k},\quad u_{k}:=u^{\mathbb{K}}_{k},\quad d_{k}:=d^{\mathbb{K}}_{k}$
for any $k\in\{0,1,\dots,\dim\mathbb{K}\}$.
@@
Let $f$ be a discrete Morse-Bott function on a finite CW complex $\mathbb{K}$. For any reduced collection $C\in\mathcal{C}_{f}$, there is a polynomial $r(t)$ with nonnegative integer coefficients such that
$\sum_{k=0}^{\dim\mathbb{K}}\#\{\sigma\in C\mid\dim\sigma=k\}t^{k}=P_{t}(C)+(1+t)r(t).$
@@
For any integer $k\in\mathbb{Z}_{\geq 0}$, the following two short exact sequences exist:
$0\rightarrow Z^{C}_{k}\rightarrow C_{k}(C)\rightarrow B_{k-1}^{C}\rightarrow 0$
@@
Then we obtain the following equalities for any integer $k\in\mathbb{Z}_{\geq 0}$:
$\operatorname{rank}C_{k}(C)=\operatorname{rank}Z^{C}_{k}+\operatorname{rank}B_{k-1}^{C}$
@@
$\sum_{k=0}^{\dim\mathbb{K}}\#\{\sigma\in C\mid\dim\sigma=k\}t^{k}-P_{t}(C)$ $=\sum_{k=0}^{\dim\mathbb{K}}\{\operatorname{rank}C_{k}(C)-b^{C}_{k}\}t^{k}$
@@
By $\operatorname{rank}B_{k-1}^{C}\geq 0$ for any integer $k\in\mathbb{Z}_{>0}$, any coefficient of $r(t)$ is nonnegative. ∎
@@
where $P_{t}(\mathbb{K})=\sum_{k}b_{k}t^{k}$ is the Poincaré polynomial of $\mathbb{K}$ and
$r(t)=\sum_{k=1}^{\dim\mathbb{K}}\bigl(\operatorname{rank}B_{k-1}-d_{k}\bigr)t^{k-1}.$
Although this proof is similar to that of [3, Theorem 3.36] in the continuous setting, to the best of the authors’ knowledge, no direct proof has been given in the literature, so we include it here for completeness.
@@
For any $k\in\mathbb{Z}_{\geq 0}$, from the two short exact sequences
@@
we obtain the following equalities:
@@
$=(1+t)\sum_{k=1}^{\dim\mathbb{K}}(\operatorname{rank}B_{k-1}-d_{k})t^{k-1}.$
@@
For a subcomplex $J$ of $\mathbb{K}$, for any $k\in\mathbb{Z}_{\geq 0}$, denote by $\partial_{J,k}:C_{k}(J)\to C_{k-1}(J)$ the boundary homomorphism. For any $k\in\mathbb{Z}_{>0}$, since $\operatorname{rank}B_{k-1}-d_{k}=u_{k}+m_{k}-\operatorname{rank}Z_{k}$, it suffices to show the inequality
(2) $u_{k}^{J}+m_{k}^{J}-\operatorname{rank}\ker(\partial_{J,k})\;\geq\;0$
@@
$f(\sigma)=\max\{f(\alpha)\mid\alpha\text{ is a maximal-dimensional cell in }K\}.$
[…]
$\begin{split}u_{k}^{K_{\sigma}}+m_{k}^{K_{\sigma}}-\operatorname{rank}\ker(\partial_{K_{\sigma},k})\geq 0\end{split}$
Because $K_{\sigma}$ is obtained from $K$ by removing the single cell $\sigma$, we obtain the following relations:
$\begin{split}u_{p}^{K}+m_{p}^{K}-(u_{p}^{K_{\sigma}}+m_{p}^{K_{\sigma}})\in\{0,1\}\\ \operatorname{rank}\ker\partial_{K,p}-\operatorname{rank}\ker\partial_{K_{\sigma},p}\in\{0,1\}\end{split}$
(3) $\operatorname{rank}\ker\partial_{K,k}=\operatorname{rank}\ker\partial_{K_{\sigma},k}$
[…]
$u_{p}^{K}+m_{p}^{K}-\operatorname{rank}\ker(\partial_{K,p})=-1$
$u_{p}^{K_{\sigma}}+m_{p}^{K_{\sigma}}-\operatorname{rank}\ker(\partial_{K_{\sigma},p})=0$
Therefore, the following equalities hold:
@@
$x:=\sum_{\alpha^{p}\in K}a_{\alpha}\alpha\in\ker\partial_{K,p},\qquad a_{\alpha}\in\mathbb{Z},$
@@
Let $f$ be a discrete Morse-Bott function on a finite CW complex $\mathbb{K}$, $k\in\{0,1,\dots,\dim\mathbb{K}\}$ an integer, and $x\in Z_{k}\setminus\{0\}$. Write
@@
the sum
$x^{\alpha}:=\sum_{\sigma\in A\cap\mathcal{C}_{f}(\alpha)}a_{\sigma}\sigma$
satisfies $x^{\alpha}\in Z^{\mathcal{C}_{f}(\alpha)}_{k}\setminus\{0\}$.
@@
Put $C:=\mathcal{C}_{f}(\alpha)$. First, we show the following statement.
@@
$\{\nu\prec\sigma\mid\nu\in C\}=\emptyset$ for any cell $\sigma\in A\setminus C$.
@@
Fix any cell $\sigma\in A\setminus C$. Define $N_{\sigma}:=\{\nu\prec\sigma\mid\nu\in C\}$. Assume that $N_{\sigma}\neq\emptyset$. Fix any cell $\nu\in N_{\sigma}$. Since $\nu\in C$, the cell $\nu$ is not upward noncritical and so
@@
$\{\nu\in\mathbb{K}\mid\sigma\in A,\;\nu\prec\sigma\}$ $=\{\nu\in\mathbb{K}\mid\sigma\in A,\;\nu\prec\sigma,\;\nu\notin C\}$
$\quad\cup\{\nu\in\mathbb{K}\mid\sigma\in A\cap C,\;\nu\prec\sigma,\;\nu\in C\}.$
@@
$=\sum_{\sigma\in A\cap C}\sum_{\begin{subarray}{c}\nu\prec\sigma\\ \nu\in C\end{subarray}}a_{\sigma}[\sigma:\nu]\nu+\sum_{\sigma\in A}\sum_{\begin{subarray}{c}\nu\prec\sigma\\ \nu\notin C\end{subarray}}a_{\sigma}[\sigma:\nu]\nu$
@@
$0$ $=\sum_{\sigma\in A\cap C}\sum_{\begin{subarray}{c}\nu\prec\sigma\\ \nu\in C\end{subarray}}a_{\sigma}[\sigma:\nu]\nu=\partial^{C}_{k}x^{\alpha}$
@@
Let $f:\mathbb{K}\to\mathbb{R}$ be a function, $k\in\{0,1,\dots,\dim\mathbb{K}\}$ an integer, and $J$ a subcomplex of $\mathbb{K}$. Then $\tilde{m}^{J}_{k},\tilde{u}^{J}_{k},\tilde{d}^{J}_{k}$ are defined as follows:
$\tilde{m}^{J}_{k}$ $:=\#\{\text{weakly critical $k$-cells for $f|_{J}$}\}$
$\tilde{u}^{J}_{k}$ $:=\sum_{\sigma^{k}\in J}\#\{\tau\in J\mid\sigma\overset{\mathrm{snc}}{\prec}\tau\}$
$\tilde{d}^{J}_{k}$ $:=\sum_{\sigma^{k}\in J}\#\{\nu\in J\mid\nu\overset{\mathrm{snc}}{\prec}\sigma\}$
If $f$ is Morse-Bott, then the following relations hold:
$c^{J}_{k}$ $=\tilde{m}^{J}_{k}+\tilde{u}^{J}_{k}+\tilde{d}^{J}_{k}\hskip 20.0pt\text{ for any $k\in\{0,1,\dots,\dim\mathbb{K}\}$}$
$\tilde{u}^{J}_{k}$ $=\tilde{d}^{J}_{k+1}\hskip 61.0pt\text{ for any $k\in\{0,1,\dots,\dim\mathbb{K}-1\}$}$
@@
Let $f$ be a discrete Morse-Bott function on a finite CW complex $\mathbb{K}$. Then there is a polynomial $R(t)$ with nonnegative integer coefficients such that
(7) $\sum_{C\in\mathcal{C}_{f}}P_{t}(C)\;=\;P_{t}(\mathbb{K})+(1+t)R(t)$
@@
The relation $\nabla_{s}f=-\nabla g$ implies $d^{g}_{k}=\tilde{d}_{k}$ and $u^{g}_{k}=\tilde{u}_{k}$. Then $m^{g}_{k}=\tilde{m}_{k}$. Therefore, the following holds:
@@
$\implies$ $\sum_{k=0}^{\dim\mathbb{K}}\sum_{C\in\mathcal{C}_{f}}\#\{\sigma\in C\mid\dim\sigma=k\}t^{k}=P_{t}(\mathbb{K})+(1+t)r(t)$
$\implies$ $\sum_{C\in\mathcal{C}_{f}}(P_{t}(C)+(1+t)\sum_{k=1}^{\dim\mathbb{K}}(\operatorname{rank}B_{k-1}^{C})t^{k-1})=P_{t}(\mathbb{K})+(1+t)r(t)$
$\implies$ $\sum_{C\in\mathcal{C}_{f}}P_{t}(C)=P_{t}(\mathbb{K})+(1+t)$
$\hskip 110.0pt\sum_{k=1}^{\dim\mathbb{K}}(\operatorname{rank}B_{k-1}-d^{g}_{k}-\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}B_{k-1}^{C})t^{k-1}.$
@@
$\operatorname{rank}C_{k}(C)$ $=\operatorname{rank}Z^{C}_{k}+\operatorname{rank}B_{k-1}^{C}$
$c^{J}_{k}$ $=\tilde{m}^{J}_{k}+\tilde{u}^{J}_{k}+\tilde{d}^{J}_{k}$
@@
$\operatorname{rank}B_{k-1}-\tilde{d}_{k}-\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}B_{k-1}^{C}=\tilde{u}_{k}+\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}Z^{C}_{k}-\operatorname{rank}Z_{k}.$
For any subcomplex $J$ of $\mathbb{K}$, put $\mathcal{C}_{J}:=\mathcal{C}_{f\mid_{J}}$. Then it suffices to show the following inequality:
(8) $\tilde{u}^{J}_{k}+\sum_{C\in\mathcal{C}_{J}}\operatorname{rank}Z^{C}_{k}-\operatorname{rank}\ker\partial_{J,k}\geq 0.$
@@
$M:=\{\sigma\in K\mid\sigma\text{ is maximal-dimensional in }K\}.$
@@
By definition of regularity of faces, we obtain that $P_{t}(C)=0$ for any reduced collection $C\in\mathcal{C}_{f}$ with $C=\{\sigma,\tau\}$ with $\sigma^{p}\overset{\mathrm{reg}}{\prec}\tau$. Suppose that $f$ is a discrete Morse function. By Corollary 3.3, for any reduced collection $C\in\mathcal{C}_{f}$, we obtain that $C\leq 2$ and $C=L(C)$.
@@
For any reduced collection $C\in\mathcal{C}_{f}$ with #C=2\#C=2, we have that $P_{t}(C)=0$. For any reduced collection $C\in\mathcal{C}_{f}$ with #C=1\#C=1, by Remark 4, #C=#L(C)=1\#C=\#L(C)=1 implies $C=L(C)=\{\text{a critical cell of }f|_{C}\}$ and so $P_{t}(C)=t^{\dim C}$. Therefore, the desired equality holds. ∎
@@
For any reduced collection $C\in\mathcal{C}_{f}$ with #C=1\#C=1, the definition of $B_{k-1}^{C}$ implies $\operatorname{rank}B_{k-1}^{C}=0$. For any reduced collection $C\in\mathcal{C}_{f}$ with #C=2\#C=2, we have $C=\{\nu\prec\sigma^{p}\}$ and so $\operatorname{rank}B_{p-1}^{C}=\operatorname{rank}\operatorname{im}\partial_{p}^{C}=1$. By definition of $d_{k}$, we have the following equalities:
$d_{k}$ $=\sum_{\sigma^{k}\in\mathbb{K}}\#\{\nu\in\mathbb{K}\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}$
$=\sum_{\sigma^{k}\in\mathbb{K}}\#\{\nu\prec\sigma\mid f(\nu)>f(\sigma)\}+\sum_{\sigma^{k}\in\mathbb{K}}\#\{\nu\prec\sigma\mid f(\nu)=f(\sigma)\}$
$=\tilde{d_{k}}+\#\{C\in\mathcal{C}_{f}\mid C=\{\nu\prec\sigma^{k}\},~\nu,\sigma\in\mathbb{K}\}$
$=\tilde{d_{k}}+\sum_{C\in\mathcal{C}_{f}}\operatorname{rank}B_{k-1}^{C}.$
In the sequence of equalities above, by $C\leq 2$, the third equality holds. ∎
@@
Acknowledgements
The authors would like to thank Professor Jesus Gonzalez for his helpful comments.